OpenStudy (anonymous):

limit X approaches 0 1-cosx / x

OpenStudy (anonymous):

you will have to apply La Hospital Rule

OpenStudy (dan815):

lol

OpenStudy (anonymous):

actually you don't what are you allowed to use for this?

OpenStudy (anonymous):

my head

OpenStudy (anonymous):

this become 0/0

OpenStudy (dan815):

hahaha

OpenStudy (anonymous):

do you mean (1-cos x)/x or 1-(cosx)/x

OpenStudy (anonymous):

Evaluate each of the following limits or show that they do not exist. Show Your work! limit x approach 0 ....(1-cosx)/(x) Trick question? just zero?

OpenStudy (anonymous):

one way is to note that this is the negative of the derivative of cosine at \(x=0\)

OpenStudy (anonymous):

another way is to recall in from your calc book that \[\lim_{x\to 0}\frac{\cos(x)-1}{x}=0\]

OpenStudy (anonymous):

it probably uses something called the "squeeze theorem" to prove it

OpenStudy (anonymous):

it's lim sinx approach 0

OpenStudy (anonymous):

Yea but we are not using infinitive limits

OpenStudy (anonymous):

Unless they want me to use them to evaluate and prove? Idk its my last limit problem on my practice final and i just want to put 1-1 / 0 = 0/0 but i think that would be wrong

OpenStudy (anonymous):

So... DNE i suppose

OpenStudy (dan815):

maann just use l'hospital rules why get so complicated xD

OpenStudy (anonymous):

http://www.youtube.com/watch?v=vNtJ87U8G3I that can be helpful

OpenStudy (dan815):

haha well thats convenient

OpenStudy (anonymous):

i don't want to break the rules ,and i want to help so.......^_^

OpenStudy (anonymous):

whats breaking the rules?

OpenStudy (anonymous):

the forum code

OpenStudy (dan815):

http://www.youtube.com/watch?v=igJdDN-DPgA watch this too its the squeeze theorm on how sin x / x = 1

OpenStudy (anonymous):

it is not complicated at all l'hopitals rule for this is inappropriate because i am going to guess that this comes way before l'hopital, and maybe even before differentiation

OpenStudy (anonymous):

the limit is 0, not "DNE"

OpenStudy (anonymous):

one easy way to find it would be to graph \[y=\frac{1-\cos(x)}{x}\] and you would see that at \(0\) it is zero

OpenStudy (anonymous):

Haha I see, I think l'hopital would work though i recognize learning it at one point, just dont remmeber it quite well, i remmeber there was a lot of situations it didnt work so i threw it out the window i think